OpenStudy (anonymous):

Which of the following is a solution to the equation log 4 x + log 4 (x – 3) = 1 ? i got x=6/5 am i right??

OpenStudy (anonymous):

@seesawn

OpenStudy (anonymous):

@Blacksteel

OpenStudy (anonymous):

@DarkFairyTale

OpenStudy (anonymous):

@jh

OpenStudy (tkhunny):

How did you manage that result?

OpenStudy (anonymous):

i have no idea

OpenStudy (tkhunny):

Well, if you've truly no idea, then your answer is not correct. Please demonstrate your knowledge of logarithm properties.

OpenStudy (anonymous):

my choices are confusing a.6/5 b.2 c.-3 d.4 i know that log properties are where you graph a linear function and it is the inverse of an exponential function

OpenStudy (tkhunny):

We're not really concerned about function definitions on this one. The task is to solve the equation. Here's a property: log(a) + log(b) = log(a*b) You can use this property to begin your voyage toward a solution.

OpenStudy (anonymous):

so i plug in the numbers for the eqquation?

OpenStudy (dumbcow):

haha that was a bit misleading...its multiple choice and you said it as if you were checking an answer when really you just guessed option A ...lol

OpenStudy (anonymous):

might as well be honest and say i did guess because i was clueless on this one

OpenStudy (tkhunny):

There is no such thing as "plug in". I suppose you could rely on checking given answers by simply substituting them, but this will not help you learn to solve such problems completely. Use that property I gave you and rewrite the original equation.

OpenStudy (dumbcow):

they need to stop with multiple choice so students learn to actually apply math skills and not just eliminate wrong answers by guessing and checking :{

OpenStudy (anonymous):

yeah my dad says the same thing

OpenStudy (dumbcow):

it makes it easier to grade though...hence the abundance of standardized tests...ughh

OpenStudy (tkhunny):

Will we be chatting or will the OP use the log properties?

OpenStudy (dumbcow):

anyway, use this property from @tkhunny Here's a property: log(a) + log(b) = log(a*b)

OpenStudy (anonymous):

oh i just remembered how to do this the logs cancel out because they are the same and i just solve (x-3)=1, so i just add 3 tyo both sides and i get 4 for my answer?

OpenStudy (tkhunny):

I hope you do not remember that, because it makes no sense.

OpenStudy (dumbcow):

haha @cowgirlworld what you are thinking of is: \[\log_{4}(x-3) = \log_{4} 1 \] the logs cancel because same base and they are set equal to each other not the case for posted question though

OpenStudy (anonymous):

but is 4 the correct answer

OpenStudy (tkhunny):

Please stop guessing. Use the property: log(a) + log(b) = log(a*b). Show your algebra skills, not your dart-throwing skills.